OpenStudy (anonymous):

Given that f(x) = x2 + 2x + 3 and g(x) = quantity of x plus four, over three, solve for f(g(x)) when x = 2

OpenStudy (anonymous):

\[f(x) = x^{2}+2 \times x + 3 ; g(x)= (x+4)^{3}\]

OpenStudy (anonymous):

am I right till now?

OpenStudy (anonymous):

f=11 g=216 f(g(x))= 4752